OpenStudy (anonymous):

Name the property illustrated by each statement: If h + 12 = 22, then h = 10.

OpenStudy (lgbasallote):

hint: h + 12 = 22 h + 12 - 12 = 22 - 12 h = 10 does that help?

OpenStudy (anonymous):

Not really, What property is it?

OpenStudy (lgbasallote):

what did i do in the middle step?

OpenStudy (anonymous):

Is it Reflexive Property?

OpenStudy (lgbasallote):

answer my question...what did i do in the middle step?

OpenStudy (lgbasallote):

h+12 - 12 = 22 -12 what did i do there?

OpenStudy (anonymous):

You added and subtracted

OpenStudy (lgbasallote):

remember..the original equation was h + 12 = 22 then i SUBTRACTED both sides by 12

OpenStudy (lgbasallote):

does that give you a hint which proeprty it is?

OpenStudy (anonymous):

Would it be Transitive?